### v1

# QVGen: Pushing the Limit of Quantized Video Generative Models

Abstract
Video diffusion models (DMs) have enabled high-quality video synthesis. Yet, their substantial computational and memory demands pose serious challenges to real-world deployment, even on high-end GPUs. As a commonly adopted solution, quantization has achieved notable successes in reducing cost for image DMs, while its direct application to video DMs remains ineffective. In this paper, we present QVGen, a novel quantization-aware training (QAT) framework tailored for high-performance and inference-efficient video DMs under extremely low-bit quantization (i.e., 4-bit or below). We begin with a theoretical analysis demonstrating that reducing the gradient norm is essential to facilitate convergence for QAT. To this end, we introduce auxiliary modules ($\Phi$) to mitigate large quantization errors, leading to significantly enhanced convergence. To eliminate the inference overhead of $\Phi$, we propose a rank-decay strategy that progressively eliminates $\Phi$. Specifically, we repeatedly employ singular value decomposition (SVD) and a proposed rank-based regularization ${\bm{\gamma}}$ to identify and decay low-contributing components. This strategy retains performance while zeroing out inference overhead. Extensive experiments across $4$ state-of-the-art (SOTA) video DMs, with parameter sizes ranging from $1.3$B14similar-toabsent14\,$\sim$ 14∼ 14B, show that QVGen is the first to reach full-precision comparable quality under 4-bit settings. Moreover, it significantly outperforms existing methods. For instance, our 3-bit CogVideoX-2B achieves improvements of $+25.28$ in Dynamic Degree and $+8.43$ in Scene Consistency on VBench.
1 Introduction
Recently, advancements in artificial intelligence-generated content (AIGC) have led to significant breakthroughs in text [60, 7], image [68, 31], and video synthesis [64, 28]. The development of video generative models, driven by the powerful diffusion transformer (DiT) [50] architecture, has been particularly notable. Leading video diffusion models (DMs), such as closed-source OpenAI Sora [49] and Kling [29], and open-source Wan  [63] and CogVideoX [72], can successfully model motion dynamics, semantic scenes, physical consistency, etc. Despite their impressive performance, these models demand high computational resources and substantial peak memory, especially when generating long videos at a high resolution. For example, Wan 14B [64] requires more than 30 minutes and 50GB of GPU memory to generate a 10-second $720p$ resolution video clip on a single $H100$ GPU.
(a) BF16
(b) W4A4 QVGen (Ours)
(c) W4A4 EfficientDM [17]
(d) W4A4 Q-DM [35]
(e) W4A4 LSQ [8]
(f) W4A6 SVDQuant [32]
Text prompt: “In the haunting backdrop of a war-torn city, where ruins and crumbled walls tell a story of devastation, a poignant close-up frames a young girl. Her face is smudged with ash, a silent testament to the chaos around her. Her eyes glistening with a mix of sorrow and resilience, capturing the raw emotion of a world that has lost its innocence to the ravages of conflict.”
[…]
Even worse, deploying such models is infeasible on most customer-grade PCs, let alone resource-constrained edge devices. As a result, their practical applications across various platforms face considerable challenges.
In light of these problems, model quantization, which maps high-precision (e.g., FP16/BF16) data to low-precision (e.g., INT8/INT4) formats, stands out as a compelling solution. For instance, employing 4-bit models with fast kernel implementation can achieve a significant 3×3$\times$3 × speedup ratio with about 4×4$\times$4 × model size reduction compared with floating-point models on NVIDIA RTX4090 GPUs [32]. However, quantizing video DMs is more challenging than quantizing image DMs, and it has not received adequate attention. As shown in Fig. 1, applying prior high-performing approaches [32, 8, 35, 17] to quantize a video DM into ultra-low bits ($\leq$ 4 bits) is ineffective. In contrast to post-training quantization (PTQ), quantization-aware training (QAT) can obtain superior performance through training quantized weights. Nevertheless, it still leads to severe video quality degradation, as demonstrated by Fig. 1 (a) vs. (c)-(e). This highlights the need for an improved QAT framework to preserve video DMs’ exceptional performance under 4-bit or lower quantization.
In this work, we present a novel QAT framework, termed QVGen. It aims to improve the convergence without additional inference costs of low-bit Quantized DMs for Video Generation.
Specifically, we first provide a theoretical analysis showing that minimizing the gradient norm $\|{\bm{g}}_{t}\|_{2}$ is the key to improving the convergence of QAT for video DMs. Motivated by this finding, we introduce auxiliary modules $\Phi$ for the quantized video DM to mitigate quantization errors. These modules effectively help narrow the discrepancy between the discrete quantized and full-precision models, leading to stable optimization and largely reduced $\|{\bm{g}}_{t}\|_{2}$. The quantized DM thus achieves better convergence. Our observation also implies that the significant performance drops (Fig. 1) of the existing SOTA QAT method [35] may result from its high $\|{\bm{g}}_{t}\|_{2}$ (Fig. 3).
Moreover, to adopt $\Phi$ for improving QAT while avoiding its substantial inference overhead, we progressively remove $\Phi$ during training. Upon further analysis, we have found that the amount of small singular values in ${\mathbf{W}}_{\Phi}$ (the weight of $\Phi$) increases throughout the training process. This indicates that the quantity of low-contributing components in ${\mathbf{W}}_{\Phi}$, which are related to small singular values [75, 71], grows during QAT. As a result, an increasing number of these components can be removed with minimal impact on training. Leveraging this insight, we introduce a rank-decay strategy to progressively shrink ${\mathbf{W}}_{\Phi}$. To be more specific, singular value decomposition (SVD) is first applied to recognize ${\mathbf{W}}_{\Phi}$’s low-impact components. Then, a rank-based regularization ${\bm{\gamma}}$ is utilized to gradually decay these components to $\varnothing$. Such processes (i.e., decompose and then decay) are repeated until ${\mathbf{W}}_{\Phi}$ is fully eliminated, which also means that $\Phi$ is removed. In terms of results, this strategy incurs minimal performance impact while getting rid of the inference overhead.
To summarize, our contributions are as follows:
  • We introduce a general-purpose QAT paradigm, called QVGen. To our knowledge, this is the first QAT method for video generation and achieves effective 3-bit and 4-bit quantization.
  • To optimize extremely low-bit QAT, we enhance a quantized DM with auxiliary modules ($\Phi$) to reduce the gradient norm. Our theoretical and empirical analysis validates the effectiveness of this method in improving convergence.
  • To eliminate the significant inference overhead introduced by $\Phi$, we propose a rank-decay strategy that progressively shrinks $\Phi$. It iteratively performs SVD and applies a rank-based regularization ${\bm{\gamma}}$ to obtain and decay low-impact components of $\Phi$, respectively. As a result, this method incurs minimal impact on performance.
  • Extensive experiments across advancing CogVideoX [72] and Wan [64] families demonstrate the SOTA performance of QVGen. Notably, our W4A4 model is the first time to show full-precision comparable performance. In addition, we apply QVGen to Wan 14B, one of the largest SOTA open-source models, and observe negligible performance drops on VBench-2.0 [78].
2 Related Work
Video diffusion models. Building upon the remarkable success of diffusion models (DMs) [19, 56, 5] in image generation [31, 68], the exploration in the field of video generation [72, 64, 28] is also becoming popular. In contrast to convolution-based diffusion models [22, 4], the success of OpenAI Sora [49] has spurred researchers to adopt the diffusion transformer (DiT) [50] architecture and scale it up for high-quality video generation. However, advanced video DiTs [64, 72, 28] often involve billions of parameters, lengthy multi-step denoising, and intensive computation over long frame sequences. This results in substantial time and memory overhead, which limits their practical deployment. To enable faster video generation, some works have introduced step-distillation [73, 39] on pre-trained models to shorten the denoising trajectory. Others focus on efficient attention [74, 67], feature caching [45, 83], or parallel inference [11, 10] to accelerate per-step computations. Moreover, to achieve memory-efficient inference, existing research has explored efficient architecture design [66, 40], structure pruning [2, 77], and model quantization [76, 6, 59]. These methods aim to achieve both model size and computational cost reduction.
Model quantization. Quantization [26] is a predominant technique for minimizing storage and accelerating inference. It can be categorized into post-training quantization (PTQ) [37, 46] and quantization-aware training (QAT) [61, 82, 8, 12]. PTQ compresses models without re-training, making it fast and data-efficient. Nevertheless, it may result in suboptimal performance, especially under ultra-low bit-width (e.g., 3/4-bit). Conversely, QAT applies quantization during training or finetuning and typically achieves higher compression rates with less performance degradation. For DMs, previous quantization research [35, 32, 17, 34, 18, 23, 24, 55, 65, 54, 63] primarily focuses on image generation. Video DMs, which incorporate complex temporal and spatial modeling, are still challenging for low-bit quantization. QVD [59] and Q-DiT [6] first apply PTQ to convolution-based [70, 13] and DiT-based [81] video DMs, respectively. Furthermore, ViDiT-[76] employs mixed-precision and fine-grained PTQ to improve performance. However, it experiences video quality loss with 8-bit activation quantization and has not been extended to more advanced models [72, 64]. Consequently, QAT for advanced video DMs is urgently needed.
In this work, we identify and address the ineffective low-bit QAT for the video DM. Our proposed method significantly enhances model performance and incurs zero inference overhead in ultra-low-bit settings. Moreover, our framework is orthogonal to existing QAT methods, which target gradient estimation [12, 8], oscillation reduction [47, 35], mixed-precision schemes [61, 51], etc.
3 Preliminaries
Video diffusion modeling. The video DM [22, 81] extends image diffusion frameworks [35, 56] into the temporal domain by learning dynamic inter-frame dependencies. Let ${\bm{x}}_{0}\in\mathbb{R}^{f\times h\times w\times c}$ be a latent video variable, where $f$ denotes the count of video frames, each of size $h\times w$ with $c$ channels. DMs are trained to denoise samples generated by adding random Gaussian noise ${\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to ${\bm{x}}_{0}$:
$\begin{array}[]{ll}{\bm{x}}_{\tau}=\alpha_{\tau}{\bm{x}}_{0}+\sigma_{\tau}{\bm{\epsilon}},\end{array}$ (1)
[…]
$\begin{array}[]{ll}\mathcal{L}(\theta)=\mathbb{E}_{{\bm{x}}_{0},{\bm{\epsilon}},\mathcal{C},\tau}[\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{\tau},\mathcal{C},\tau)\|^{2}_{F}],\end{array}$ (2)
where $\mathcal{C}$ represents conditional guidance, like texts or images, and $\|\cdot\|_{F}$ denotes the Frobenius norm. Additionally, v-prediction [53] (i.e., predict $\frac{d{\bm{x}}_{\tau}}{d\tau}$) is also a prevailing option [72, 64, 28] as the target. During inference, we can employ ${\bm{\epsilon}}_{\theta}$ with various sampling methods [43, 79] progressively denoising from a random Gaussian noise ${\bm{x}}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to a clean video variable. The raw video is obtained by decoding the variable via a video variational auto-encoder (VAE) [72, 14, 64].
Quantization. The current video DM based on the diffusion transformer (DiT) [50] architecture primarily consists of linear layers. Given an input ${\mathbf{X}}\in\mathbb{R}^{m\times k}$, a full-precision linear layer with weight ${\mathbf{W}}\in\mathbb{R}^{n\times m}$ and the layer’s quantized version can be formulated as:
$\begin{array}[]{ll}{\mathbf{Y}}={\mathbf{W}}{\mathbf{X}},\hat{{\mathbf{Y}}}={\mathcal{Q}}_{b}({\mathbf{W}}){\mathcal{Q}}_{b}({\mathbf{X}}),\end{array}$ (3)
where ${\mathbf{Y}}\in\mathbb{R}^{n\times k}$ and $\hat{{\mathbf{Y}}}\in\mathbb{R}^{n\times k}$  [^2] Here, $n$ denotes the output channel, $m$ signifies the token dimension, and $k$ is the token number. We omit the batch dimension for clarity. represent the outputs of the full-precision (e.g., FP16/BF16) and quantized linear layers, respectively. ${\mathcal{Q}}_{b}(\cdot)$ denotes the function of $b$-bit quantization. In this paper, we adopt asymmetric uniform quantization. For example, ${\mathcal{Q}}_{b}({\mathbf{X}})$ can be represented as:
$\begin{array}[]{ll}{\mathcal{Q}}_{b}({\mathbf{X}})=(\mathrm{clamp}(\lfloor\frac{{\mathbf{X}}}{s}\rceil+z,0,2^{b}-1)-z)\times s,\mathrm{where}~{}s=\frac{\max({\mathbf{X}})-\min({\mathbf{X}})}{2^{b}-1},z=-\lfloor\frac{\min({\mathbf{X}})}{s}\rceil.\end{array}$ (4)
[…]
$\begin{array}[]{ll}\frac{\partial{\mathcal{Q}}_{b}({\mathbf{X}})}{\partial{\mathbf{X}}}={\mathbb{I}}_{0\leq\lfloor\frac{{\mathbf{X}}}{s}\rceil+z\leq 2^{b}-1}.\end{array}$ (5)
Similar to existing works [35, 80], we employ the full-precision model as the teacher to guide the training of the quantized model in a knowledge distillation-based (KD-based) manner. Therefore, the training loss can be defined as:
$\mathcal{L}=\mathbb{E}_{{\bm{x}}_{0},\mathcal{C},\tau}[\|\hat{\bm{\epsilon}}_{\theta}({\bm{x}}_{\tau},\mathcal{C},\tau)-{\bm{\epsilon}}_{\theta}({\bm{x}}_{\tau},\mathcal{C},\tau)\|_{F}^{2}],$ (6)
where $\hat{\bm{\epsilon}}_{\theta}$ denotes the quantized denoiser of a video DM.
4 QVGen
Figure 2: Overview of the proposed QVGen. (a) This framework integrates auxiliary modules $\Phi$ to improve training convergence (Sec. 4.1). (b) To maintain performance while eliminating inference overhead induced by $\Phi$, we design a rank-decay schedule that progressively shrinks the entire $\Phi$ to $\varnothing$ through iteratively applying the following two strategies (Sec. 4.2): 1) SVD to identify the low-impact components in $\Phi$; 2) A rank-based regularization ${\bm{\gamma}}$ to decay the identified components to $\varnothing$.
Considering the substantial video-quality drops (see Fig. 1) observed in existing QAT methods [35, 17, 8], we believe that the quantized video DM suffers from poor convergence. In the following subsections, we propose QVGen (see Fig. 2) to address this issue while maintaining inference efficiency.
4.1 Improving Convergence with Auxiliary Modules $\Phi$
To begin with, we analyze the convergence of a quantized video DM using the regret, which is widely used in the analysis of deep learning optimizers [27, 44]. It is defined as:
$\begin{array}[]{ll}R(T)=\sum^{T}_{t=1}f_{t}({\bm{\theta}}_{t})-f_{t}({\bm{\theta}}^{*}),\end{array}$ (7)
where $T$ signifies the total number of training iterations and $f_{t}(\cdot)$ is the unknown cost function at iteration $t$. Here, ${\bm{\theta}}_{t}$ represents the parameters of the quantized video DM at training step $t$, constrained within a convex compact set $\mathbb{S}^{d}$, while ${\bm{\theta}}^{*}=\arg\,\min_{{\bm{\theta}}\in\mathbb{S}^{d}}\sum^{T}_{t=1}f_{t}({\bm{\theta}})$ is the optimal parameters. In QAT, ${\bm{\theta}}_{t}$ is updated by gradient descent, with the learning rate $\eta_{t}$ and gradient ${\bm{g}}_{t}$, as:
$\begin{array}[]{ll}{\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}{\bm{g}}_{t}.\end{array}$ (8)
Theorem 4.1.
Assume that $f_{t}$ is convex [^3] This may not hold for deep networks. However, it facilitates the analysis of deep learning models [27, 44]. and $\forall{\bm{\theta}}_{i},{\bm{\theta}}_{j}\in\mathbb{S},\|{\bm{\theta}}_{i}-{\bm{\theta}}_{j}\|_{\infty}\leq D_{\infty}$. Then the average regret is upper-bounded as:
$\begin{array}[]{ll}\frac{R(T)}{T}\leq\frac{dD_{\infty}}{2T\eta_{T}}+\frac{1}{T}\sum^{T}_{t=1}\frac{\eta_{t}}{2}\|{\bm{g}}_{t}\|_{2}^{2}.\end{array}$ (9)
A smaller value of $\frac{R(T)}{T}$ implies a closer convergence to the optimum. Thm. 4.1 (with a proof provided in the appendix) suggests that for a large $T$ (i.e., $\frac{dD_{\infty}}{2T\eta_{T}}$ becomes negligible), minimizing $\|{\bm{g}}_{t}\|_{2}$ is critical for improving convergence behavior of QAT.
A lower $\|{\bm{g}}_{t}\|_{2}$ is typically observed in more stable training processes [58, 69]. Therefore, to reduce $\|{\bm{g}}_{t}\|_{2}$, we aim to stabilize the QAT process by mitigating aggressive training losses (e.g., loss spikes) [30, 33]. Specifically, we introduce a learnable auxiliary module $\Phi$ to enhance each quantized linear layer of a video DM. This module is trained to mitigate severe quantization-induced errors during QAT, thereby preventing aggressive training losses. The forward computation of such a $\Phi$-equipped layer becomes:
$\begin{array}[]{ll}\hat{{\mathbf{Y}}}={\mathcal{Q}}_{b}({\mathbf{W}}){\mathcal{Q}}_{b}({\mathbf{X}})+\Phi({\mathcal{Q}}_{b}({\mathbf{X}})),\end{array}$ (10)
where $\Phi({\mathcal{Q}}_{b}({\mathbf{X}}))={\mathbf{W}}_{\Phi}{\mathcal{Q}}_{b}({\mathbf{X}})$. Here, ${\mathbf{W}}_{\Phi}$ is initialized before QAT by the weight quantization error, defined as ${\mathbf{W}}-{\mathcal{Q}}_{b}({\mathbf{W}})$.
(a) CogVideoX-2B [72]
(b) Wan 1.3B [64]
Figure 3: (Upper) $\|{\bm{g}}_{t}\|_{2}$ vs. #steps and (Lower) training loss vs. #steps across different video DMs and 4-bit QAT methods under the same training settings.“$\Phi$” denotes our approach in Sec. 4.1.
To validate the effectiveness of $\Phi$, we conduct experiments for CogVideoX 2B [72] and Wan 1.3B [64]. Compared with the previous SOTA QAT method Q-DM [35], the proposed approach exhibits consistently lower $\|{\bm{g}}_{t}\|_{2}$ and reduced training loss, as depicted in Fig. 3. This aligns well with both the theoretical and the empirical analysis discussed earlier. Therefore, incorporating $\Phi$ in QAT effectively reduces the gradient norm and leads to better convergence for QAT. In addition, as evidenced by Fig. 3, the substantial performance degradation of Q-DM (e.g., depicted in Fig. 1) for the video generation task could be attributed to its relatively large $\|{\bm{g}}_{t}\|_{2}$.
4.2 Progressively Shrinking $\Phi$ via Rank-Decay
However, during inference, the auxiliary module $\Phi$ introduces non-negligible overhead. Concretely, $\Phi$ incurs additional matrix multiplications between $b$-bit activations ${\mathcal{Q}}_{b}({\mathbf{X}})$ and full-precision weights ${\mathbf{W}}_{\Phi}$. This is inapplicable to low-bit multiplication kernels and thus hinders inference acceleration. In addition, the storage of full-precision ${\mathbf{W}}_{\Phi}$ for each $\Phi$ leads to significant memory overhead, exceeding that of the quantized diffusion model by severalfold.
To improve QAT while eliminating the inference overhead, we propose to progressively remove $\Phi$ throughout the training process. This allows the model to benefit from $\Phi$ during QAT, while ultimately yielding a standard quantized model [35, 8] with no extra inference cost. To achieve this goal, a straightforward solution is to decay all parameters of $\Phi$ directly. However, we have noticed that it is ineffective and suboptimal (see Tab. 5.3). This calls for a fine-grained decay strategy.
(a) CogVideoX-2B [72]
(b) Wan 1.3B [64]
Figure 4: Singular value variation in ${\mathbf{W}}_{\Phi}$ across training iterations for 4-bit video DMs. We visualize the average of the singular values $\{\sigma_{s}\}_{s=1,2,\ldots,2^{10}}\cup\{\sigma_{d}\}$ across layers of all Attention blocks [62] and feed-forward networks (FFNs), respectively. “00 step” denotes the initialization state before the start of QAT.
Based on the above analysis, we investigate the contribution of fine-grained components in $\Phi$. Specifically, we apply singular value decomposition (SVD) [32, 36] to ${\mathbf{W}}_{\Phi}$ at various training steps:
$\begin{array}[]{ll}{\mathbf{W}}_{\Phi}=\sum_{s=1}^{d}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{T},\end{array}$ (11)
where $d=\min\{n,m\}$ and $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{d}$ are the singular values. The vectors ${\bm{u}}_{s}$ and ${\bm{v}}_{s}$ denote the left and right singular vectors associated with $\sigma_{s}$, respectively. By tracking the evolution of the average $\sigma_{s}$, we observe two key findings (exemplified by Fig. 4 (a) Attention):
[…]
These findings suggest that an increasing number of orthonormal directions $\{{\bm{u}}_{s},{\bm{v}}_{s}\}$ contribute little, as their associated singular values $\sigma_{s}$ are small [75, 71]. Hence, as training proceeds, only an increasingly low-rank portion of $\Phi$ (i.e., $\Sigma^{d^{\prime}}_{s=1}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{T}$, where $d^{\prime}<d$) is needed, and the remaining components can be decayed without noticeably affecting performance. Motivated by this, we propose a novel rank-decay schedule that progressively shrinks $\Phi$ by repeatedly identifying and eliminating the above-mentioned low-impact parts. First, to attain these parts, we reformulate the computation of $\Phi$ as:
$\begin{array}[]{ll}\Phi({\mathcal{Q}}_{b}({\mathbf{X}}))={\mathbf{L}}{\mathbf{R}}{\mathcal{Q}}_{b}({\mathbf{X}}),\end{array}$ (12)
where ${\mathbf{L}}=[\sqrt{\sigma_{1}}{\bm{u}}_{1},\ldots,\sqrt{\sigma_{r}}{\bm{u}}_{r}]\in\mathbb{R}^{n\times r}$ and ${\mathbf{R}}=[\sqrt{\sigma_{1}}{\bm{v}}_{1},\ldots,\sqrt{\sigma_{r}}{\bm{v}}_{r}]^{T}\in\mathbb{R}^{r\times m}$ for a given rank $r$. In practice, we set $r\ll d$ to reduce training costs, as ${\mathbf{W}}_{\Phi}$ already exhibits a non-negligible number of small singular values before QAT. Consequently, $\sqrt{\sigma_{s}}{\bm{u}}_{s}$ and $\sqrt{\sigma_{s}}{\bm{v}}_{s}$ are the components in ${\mathbf{W}}_{\Phi}$ represent the $s$-th level of contribution. Then, with a rank-based regularization ${\bm{\gamma}}$ applied, the forward computation of a quantized linear layer during training is modified as:
$\begin{array}[]{ll}\hat{{\mathbf{Y}}}={\mathcal{Q}}_{b}({\mathbf{W}}){\mathcal{Q}}_{b}({\mathbf{X}})+({\bm{\gamma}}\odot{\mathbf{L}}){\mathbf{R}}{\mathcal{Q}}_{b}({\mathbf{X}}),\end{array}$ (13)
where ${\bm{\gamma}}$ is defined as:
$\begin{array}[]{ll}{\bm{\gamma}}=\mathrm{concat}([1]_{n\times(1-\lambda)r},[u]_{n\times\lambda r})\in\mathbb{R}^{n\times r}.\end{array}$ (14)
Here, $u$ follows a cosine annealing schedule that decays from $1$ to 00, $\lambda\in(0,1]$ represents the shrinking ratio, and $\odot$ denotes element-wise multiplication. Eq. (13) and Eq. (14) allow us to progressively eliminate the low-impact components of $\Phi$ (i.e., $[\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{u}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{u}}_{r}]$ and $[\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{v}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{v}}_{r}]^{T}$). Once $u$ reaches 00, we truncate $\{{\mathbf{L}},{\mathbf{R}}\}$ to $\{{\mathbf{L}}^{\prime},{\mathbf{R}}^{\prime}\}$, and rewrite ${\mathbf{W}}_{\Phi}$ as:
$\begin{array}[]{ll}\begin{cases}{\mathbf{L}}^{\prime}={\mathbf{L}}[:,:(1-\lambda)r]\\ {\mathbf{R}}^{\prime}={\mathbf{R}}[:(1-\lambda)r,:]\end{cases}\Rightarrow{\mathbf{W}}_{\Phi}={\mathbf{L}}^{\prime}{\mathbf{R}}^{\prime}.\end{array}$ (15)
In Eq. (15), the rank of ${\mathbf{W}}_{\Phi}$ is shrunk from $r$ to $(1-\lambda)r$. During the subsequent training phase, the above procedures (i.e., both decomposition and decay) are iteratively applied. Ultimately, we fully eliminate $\Phi$ by reducing $r$ to 00, which incurs negligible impact on model performance (see Tab. 5.3). It is worth noting that we set $\lambda=\frac{1}{2}$ for Eq. (14) in this work, based on its effectiveness demonstrated in Tab. 5.3. Overall, the overview of the rank-decay schedule is exhibited in Fig. 2 (b).
5 Experiments
5.1 Implementation Details
Models. We conduct experiments on open-source SOTA video DMs, including CogVideoX-2B and 1.5-5B [72], and Wan 1.3B and 14B [64]. Classifier-free guidance (CFG[21] is used for all models, and the frame number of generated videos is fixed to $49$ for CogVideoX-2B and $81$ for the others.
Baselines. We adopt previous powerful PTQ and QAT methods as baselines: ViDiT-Q [76], SVDQuant [32], LSQ [8], Q-DM [35], and EfficientDM [17]. Since these methods were designed for image DMs or convolutional neural networks (CNNs), we adapt these works to video DMs using their open-source code (if available) or the implementation details provided in the corresponding papers. Without specific clarification, static per-channel weight quantization with dynamic per-token activation quantization, a common practice in the community [76, 41], is used for all the linear layers.
Training. We employ $16K$ captioned videos from OpenVidHQ-4M [48] as the training dataset. The AdamW [42] optimizer is utilized with a weight decay of $10^{-4}$. We employ a cosine annealing schedule to adjust the learning rate over training. During QAT, we train Wan 14B [64] and CogVideoX1.5-5B [72] for $16$ epochs on $10032\times H$ 100 GPUs and $10016\times H$ 100 GPUs, respectively. For the other DMs, we employ $8$ training epochs on $1008\times H$ 100 GPUs. Additionally, we allocate the same training iterations for each decay phase (i.e., shrinking the remaining $r$ to $(1-\lambda)r$). The same training settings are applied to all QAT baselines.
Evaluation. We select $8$ dimensions in VBench [25] with unaugmented prompts to comprehensively evaluate the performance following previous studies [76, 52]. Moreover, for quantized huge ($\geq$ 5B parameters) DMs, we additionally report the results on VBench-2.0 [78] with augmented prompts to measure the adherence of videos to physical laws, commonsense reasoning, etc.
More detailed experimental setups can be found in the appendix.
Table 1: Performance comparison across different quantization methods on VBench [25]. “†” indicates PTQ methods and “*” signifies QAT methods. “Full Prec.” denotes the BF16 model. $\clubsuit$” represents that we apply fine-grained per-group weight-activation quantization with a group size of 64 and keep some linear layers unquantized, which is the same as the official settings of SVDQuant [32] (details can be found in the appendix). Best and second-best results are highlighted in bold and underline formats, respectively. The gaps between them are marked in red subscripts.
{tabu}
[t!]l|c|llllllll Method #Bits (W/A) Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
CogVideoX-2B ($\texttt{CFG}=6.0,480p,\texttt{fps}=8$)
Full Prec. 16/16 59.15 54.49 97.43 67.78 94.79 92.82 36.24 25.06
ViDiT-Q [76] 4/6 54.72 43.01 92.18 43.22 90.76 81.02 26.25 20.41
SVDQuant [32] 4/6 58.27 47.06 95.28 40.83 92.41 87.45 27.69 21.34
SVDQuant [32]†♣ 4/4 51.60 49.40 97.69 42.22 94.03 91.78 25.67 22.89
LSQ [9] 4/4 58.73 54.20 97.57 45.00 92.97 92.41 24.06 23.17
Q-DM [35] 4/4 54.96 52.71 98.00 48.61 93.82 91.86 28.02 23.87
EfficientDM [17] 4/4 55.96 51.97 98.03 46.67 94.10 91.70 27.76 24.28
QVGen (Ours) 4/4 60.16+1.43 54.61+0.41 98.06+0.03 67.22+18.61 94.38+0.28 93.01+0.60 31.42+3.40 24.61+0.33
LSQ [9] 3/3 56.46 40.35 97.98 0.56 94.08 89.18 4.80 13.80
Q-DM [35] 3/3 50.88 40.41 98.03 5.56 93.93 87.75 7.33 15.98
EfficientDM [17] 3/3 52.86 44.58 97.13 28.61 93.15 88.26 15.42 20.42
QVGen (Ours) 3/3 58.36+1.90 50.54+5.96 98.37+0.34 53.89+25.28 94.55+0.47 90.50+1.32 23.85+8.43 22.92+2.50
Wan 1.3B ($\texttt{CFG}=5.0,480p,\texttt{fps}=16$)
Full Prec. 16/16 64.30 58.21 97.37 70.28 95.94 93.84 28.05 24.67
ViDiT-Q [76] 4/6 56.24 50.18 94.81 52.43 89.67 82.53 13.45 19.58
SVDQuant [32] 4/6 58.16 51.27 97.05 49.44 93.74 91.71 14.18 23.26
SVDQuant [32]†♣ 4/4 57.57 46.30 94.21 72.22 93.16 77.96 12.73 21.91
LSQ [9] 4/4 59.11 49.09 98.35 71.11 92.66 91.67 10.38 18.83
Q-DM [35] 4/4 60.40 52.50 97.22 76.67 93.37 89.26 13.28 21.63
EfficientDM [17] 4/4 60.70 53.57 96.18 56.39 93.74 91.70 11.77 21.19
QVGen (Ours) 4/4 63.08+2.38 54.67+1.10 98.25-0.10 77.78+1.11 94.08+0.34 92.57+0.87 15.32+2.04 23.01+1.38
LSQ [9] 3/3 58.80 46.86 98.22 23.61 91.86 89.42 0.89 15.51
Q-DM [35] 3/3 56.19 44.95 95.13 76.94 92.09 83.82 1.79 16.89
EfficientDM [17] 3/3 42.32 33.52 96.50 70.28 92.10 74.79 0.04 11.38
QVGen (Ours) 3/3 67.35+8.55 49.71+2.85 98.93+0.71 84.14+7.20 93.62+1.52 92.25+2.83 5.71+3.92 20.11+3.22
5.2 Performance Analysis
Comparison with baselines. We report VBench [25] score comparisons in Tab. 5.1. In W4A4 quantization, recent QAT methods [8, 17, 35] show non-negligible performance degradation. With W3A3, performance drops become more pronounced. In contrast, the proposed QVGen achieves substantial performance recovery in 3-bit models and comparable results to full-precision models in 4-bit quantization. Specifically, it shows higher scores or less than a $2\%$ decrease in all metrics for W4A4 CogVideoX-2B [72], except Scene Consistency. For PTQ baselines [76, 32], all fail to generate meaningful content in W4A4 per-channel and per-token settings. Therefore, we apply W4A6 quantization for these methods and also utilize fine-grained per-group W4A4 quantization for SVDQuant [32]. In these cases, W4A4 QVGen outperforms them by a large margin, particularly with $8.37$ and $14.61$ higher Aesthetic Quality and Subject Consistency for Wan 1.3B compared to W4A4 SVDQuant. In addition to these findings, we observe that for Wan 1.3B, Dynamic Degree recovers easily during QAT, even surpassing that of the full-precision model. However, for CogVideoX-2B, this metric significantly drops. Moreover, Scene Consistency is the most challenging metric to maintain across models and methods.
Beyond the quantitative results, we provide qualitative results in Fig. 1 and the appendix, where QVGen markedly improves visual quality over prior methods. Although a clear gap remains between 3-bit and 4-bit outputs, our work lays a foundation for practical 3-bit or lower video DM quantization.
Table 2: Performance for huge video DMs on VBench [25].
{tabu}
[t!]l|c|llllllll Method #Bits (W/A) Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
CogVideoX1.5-5B ($\texttt{CFG}=6.0,720p,\texttt{fps}=16$)
Full Prec. 16/16 66.25 59.49 98.42 59.72 96.57 95.28 39.14 26.18
QVGen (Ours) 4/4 66.76 59.52 98.38 64.44 95.83 94.88 28.47 24.45
QVGen (Ours) 3/3 54.44 35.85 97.23 58.89 96.48 90.17 13.27 17.15
Wan 14B ($\texttt{CFG}=5.0,720p,\texttt{fps}=16$)
Full Prec. 16/16 67.89 61.54 97.32 70.56 96.31 94.08 33.91 26.17
QVGen (Ours) 4/4 66.87 59.41 97.71 76.11 96.50 94.45 19.84 25.70
QVGen (Ours) 3/3 48.70 29.73 99.05 93.33 97.34 94.71 2.81 13.97
(c) CogVideoX1.5-5B [72]
(d) Wan 14B [64]
Figure 5: Performance for huge video DMs on VBench-2.0 [78]. Our 4-bit models exhibit a minimal drop of $\sim 1\%$ in total score.
Results for huge DMs. To demonstrate the scalability of our method, we further test two huge video DMs, including CogVideoX1.5-5B and Wan 14B, at $720p$ resolution. As illustrated in Tab. 5.2, our 3-bit and 4-bit models follow the same pattern seen in smaller models (Tab. 5.1). However, 3-bit quantization incurs much larger drops on demanding metrics such as Scene and Overall Consistency, underscoring the challenge of pushing these larger models to 3 bits. In Fig. 5, we further assess the models with VBench-2.0 [78]; the W4A4 DMs incur only negligible overall performance loss.
5.3 Ablation Studies
Table 3: Ablation results of each component. “Naive” denotes naive QAT in a KD-based manner.
{tabu}
[t!]l|llllll Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
Naive 60.40 52.50 76.67 13.28 21.63
$+\Phi$ 63.41 54.75 77.89 15.51 22.98
$+$rank-decay  63.08-0.33 54.67-0.08 77.78-0.11 15.32-0.19 23.01+0.03
To demonstrate the effect of each design, we employ W4A4 Wan 1.3B [64] with 5 dimensions of VBench [25] for ablation studies. More ablations can be found in the appendix.
Effect of different components. We evaluate the contribution of each component in Tab. 5.3. The auxiliary module $\Phi$ (Sec. 4.1) yields substantial performance improvements across all metrics. Furthermore, the rank-decay schedule (Sec. 4.2) effectively eliminates inference overhead, while inducing less than a $0.6\%$ drop in most metrics. It even leads to a slight improvement in Overall Consistency.
Choice of the shrinking ratio $\lambda$. To determine a proper shrinking ratio $\lambda$ in Eq. (14), we conduct experiments in Tab. 5.3. By maintaining the same number of training iterations for each decay phase  [^4] We also employ a fixed total number of training epochs., a small ratio results in an excessively rapid descent of $u$ in Eq. (14) from $1$ to 00, potentially destabilizing the training process. On the other hand, a larger ratio may cause the premature removal of high-contributing components during each phase. An extreme scenario would involve a $100\%$ ratio (i.e., $\lambda=1$), in which the entire ${\mathbf{W}}_{\Phi}$ is removed in a single decay phase, leading to significant performance degradation. Thus, we adopt a moderate ratio of $\lambda=\frac{1}{2}$ in QVGen.
Table 4: Results of different shrinking ratios $\lambda$ in Eq. (14) for each decay phase. $\lambda=1$ means directly decaying the entire ${\mathbf{W}}_{\Phi}$. $\lambda=\frac{1}{2}$ is used in this work.
{tabu}
[t!]c|llllll $\quad\quad\lambda\quad\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
$1/4$ 63.02 54.23 76.84 15.18 22.85
$1/2$ 63.08+0.06 54.67+0.05 77.78-0.13 15.32+0.14 23.01+0.12
$3/4$
62.89 54.62 77.91 15.04 22.89
$1$ 61.05 52.48 76.48 13.82 21.81
Table 5: Results of different initial ranks $r$ for Eq. (15). $r=0$ represents “Naive” in Tab. 5.3. We employ $r=32$ in this work.
{tabu}
[t!]c|llllll $\quad\quad r\quad\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
0 60.40 52.50 76.67 13.28 21.63
8 62.71 54.47 74.62 14.42 22.81
16 62.99 54.62 76.58 14.84 23.00
32 63.08+0.02 54.67+0.05 77.78+1.04 15.32-0.08 23.01+0.01
64 63.06 54.30 76.74 15.40 22.92
Choice of the initial rank $r$. We further present the results for different initial ranks $r$ of ${\mathbf{W}}_{\Phi}$ in Tab. 5.3. As $r$ increases, the performance gains diminish and eventually deteriorate (i.e., at $r=64$). We attribute this trend to the same issue associated with a small shrinking ratio discussed earlier. Specifically, increasing $r$ to $2r$ introduces an additional decay phase, which shortens the training time allocated to each phase and may lead to overly rapid decay (e.g., at $r=64$).
Table 6: Results of different decay strategies. Details of each method can be found in the appendix. “Rank-based” denotes the rank-decay strategy in this work.
{tabu}
[t!]c|lllll Decay Strategy Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
Sparse-based 61.15 54.06 74.24 13.86 22.52
Res. Q.-based  61.72 54.01 72.41 14.17 22.31
Rank-based 63.08+1.36 54.67+0.61 77.78+3.54 15.32+1.15 23.01+0.49
Analysis of different fine-grained decay strategies. To further demonstrate the effectiveness of our rank-decay strategy, we evaluate alternative decay strategies in Tab. 5.3. Inspired by network pruning [16, 15], we introduce a “Sparse-based” strategy that progressively prunes the largest values in ${\mathbf{W}}_{\Phi}$ during training. Additionally, motivated by residual quantization [38], we design a “Res. Q.-based” strategy, which first quantizes ${\mathbf{W}}_{\Phi}$ into 4×4$\times$4 ×4-bit tensors with the same shape and then progressively removes them one by one. Among all these methods, the “Rank-based” strategy outperforms others across all the metrics by a large margin. This result highlights the superiority of our rank-decay schedule.
Additionally, both “Sparse-based” and “Res. Q.-based” strategies require at least 1.8×1.8$\times$1.8 × training hours at the same setups compared with our “Rank-based” approach (see the appendix).
5.4 Efficiency Discussion
(a) Model Size (GB)$\downarrow$
(b) Inference Latency (s)$\downarrow$
Figure 6: Evaluation of memory compression and inference acceleration. Blue color and yellow color denote Wan [64] 1.3B and 14B, respectively.
We report the per-step latency of W4A4 DiT components on one $A800$ GPU in Fig. 6 (b), using the same settings as Sec. 5.2. Adapted from the fast CUDA kernel implementation by Ashkboos et al. [1], W4A4 QVGen achieves 1.21×1.21$\times$1.21 × and 1.44×1.$4$ speedups for Wan 1.3B and 14B, respectively. Besides, it exhibits 4$\sim$ 4$\times$ 4 × memory savings compared to the BF16 format, as shown in Fig. 6 (a). Nevertheless, we believe that the acceleration ratio could be further improved with advanced kernel-fusion techniques, which we leave to future work. In addition, it is worth noting that our QVGen adheres to standard uniform quantization, enabling drop-in deployment with existing W4A4 kernels across various devices.
6 Conclusions and Limitations
In this work, we are the first to explore the application of quantization-aware training (QAT) in video DMs. Specifically, we provide a theoretical analysis that identifies that lowering the gradient norm is essential to improve convergence. Then, we propose an auxiliary module ($\Phi$) to achieve this. Additionally, we design a rank-decay schedule to progressively eliminate $\Phi$ for zero inference overhead with minimal impact on performance. Extensive experiments for 3-bit and 4-bit quantization validate the effectiveness of our framework, QVGen. In terms of limitations, we focus on video generation in this work. However, we believe that our methods can be generalized to more tasks, e.g., image generation and natural language processing (NLP), which we will explore in the future.